OpenStudy (anonymous):

SOMEONE PLEASE HEP ME WITH THIS PROBLEM A total cost of producing x toasters per day is given by C=1/10x^2+20x+25 euros , and the selling price of each toaster is (44-1/5x)euros.How many toasters should be produced each day in order to maximize the total profit ?

OpenStudy (deoxna):

I would subtract the selling price function from the production cost function, resulting in a function for net profit. Then you would just have to find the local maximum of this new profit function.

OpenStudy (anonymous):

so i just subtract the formula given by the euros given

OpenStudy (deoxna):

Yes, and then maximize the profit, so: Profit = 1/10x^2+20x+25 - (44-1/5x) =1/10x^2 + 101/5 x - 19

OpenStudy (anonymous):

could i put that on a calculator

OpenStudy (deoxna):

Yes, and then solve for the maximum (sorry for the delay but openstudy isn't notifying of updates). Anyway, hoped it helped!

OpenStudy (anonymous):

okaii thank you for your help

OpenStudy (deoxna):

Sure thing....